OpenStudy (anonymous):

The motion of a spring-mass system is described by y=4sin(piT-pi/4) where y is the distance in feet the equilibrium position and t is time in seconds. Find the acceleration of the mass at time t=3/4 seconds. The answer will be in pi^2 units.

OpenStudy (anonymous):

Water is discharged from a pipeline ata velocity v given by v=1482p^(1/2), where p is the pressure (in psi). If the water pressure is changing at a rate of 154 psi/second, find the acceleration (dv/dt) of the water when p=34 psi

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

take the derivative with respect to t instead of p \[\frac{ dv }{ dt }=1482*(\frac{ 1 }{ 2 })*p ^{-1/2}*\frac{ dp }{ dt }\] where p= 34 and dp/dt = 154

OpenStudy (anonymous):

...did u just give me the answer??...so all i have to do is plug in

OpenStudy (anonymous):

sadly yes because people find related rates so challenging I will typically give the derivative as a way for you to check to see if you are going down the right path

OpenStudy (anonymous):

BTW you should typically post a new question as a separate question not a comment on another question

OpenStudy (anonymous):

XO thank you! yeah, i didn't realize that I had done that...I thought I had done it seperately. I'm still new at this site.